OpenStudy (anonymous):

Someone help please! Write the equation of a line that includes the point (-4, 0) and has a slope of -3/5 in slope-intercept form. A) y = -3/5x - 12/5 B) y = -3/5x - 12 C) y = -3/5x - 5 D) y = -3/5x + 12/5

OpenStudy (anonymous):

okay what you think

OpenStudy (anonymous):

I don't know that's why I need help...

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

Well the equation is y-y1=m(x-x1). -4 is your X1, and 0 is your Y1. plug them in. And plug -3/4 into M.

OpenStudy (anonymous):

y-y1=m(x-x1)

OpenStudy (anonymous):

-3/5 MY BAD.

OpenStudy (anonymous):

as @ MarrBear said y-0=-3/5(x-(-4))

OpenStudy (anonymous):

So it would be y=-3/5(x+4). Then Distribute.

OpenStudy (anonymous):

y=-3/5x-(-3/5*-4)

OpenStudy (anonymous):

What ^ No. LOL

OpenStudy (anonymous):

it is the same........ (x-(-4))=(x+4)

OpenStudy (anonymous):

y=-3/5x-12/5

OpenStudy (anonymous):

So Y = -3/5x - 12/5 is the correct answer?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Thank you so much. If I click "best response" will take give you a medal?